OpenStudy (anonymous):

dy/dx=2/x-1/x^2 where y=2 when x=-1

OpenStudy (anonymous):

y=2ln(x)+1/x +c put x=-1 get c

OpenStudy (anonymous):

\[-1=e^{(i (2n+1)\pi)}\]\[\ln(-1)=i(2n+1)\pi\] 2=2i(2n+1) pi -1+c c=3-2i(2n+1)pi \[y=2\ln(x)+\frac1x+3-2i(2n+1)\pi\] where\[n=(.....,-1,0,1,2,3...)\]

OpenStudy (anonymous):

did u get that?????

OpenStudy (anonymous):

yes but the book says the answer should be y=ln(x)^2+(1/x)-2

OpenStudy (anonymous):

2ln(x)=ln(x)^2....and is the point x=-1 given in your book???

OpenStudy (anonymous):

my mistake you are right thank you